find three consecutive even integers such that the largest is three times the smallest.: find three consecutive even integers such that the largest is three times the smallest.
Answer by orca(336)

Let the three consecutive even numbers be n-2,n,n+2.
As the largest is three times the smallest, we have:
n+2=3(n-2)
Solving for n, we obtain:
n+2=3n-6
8=2n
8/2=n
4=n
So n = 4
Thus the three consecutive even numbers is 2, 4, 6.